# A Systematization of the Wagner Framework: Graph Theory Conjectures and Reinforcement Learning

### v1

@@ Abstract
In 2021, Adam Zsolt Wagner proposed an approach to disprove conjectures in graph theory using Reinforcement Learning (RL). Wagner’s idea can be framed as follows: consider a conjecture, such as a certain quantity $f(G)<0$ for every graph $G$; one can then play a single-player graph-building game, where at each turn the player decides whether to add an edge or not. The game ends when all edges have been considered, resulting in a certain graph $G_{T}$, and $f(G_{T})$ is the final score of the game; RL is then used to maximize this score. This brilliant idea is as simple as innovative, and it lends itself to systematic generalization. Several different single-player graph-building games can be employed, along with various RL algorithms. Moreover, RL maximizes the cumulative reward, allowing for step-by-step rewards instead of a single final score, provided the final cumulative reward represents the quantity of interest $f(G_{T})$. In this paper, we discuss these and various other choices that can be significant in Wagner’s framework. As a contribution to this systematization, we present four distinct single-player graph-building games. Each game employs both a step-by-step reward system and a single final score. We also propose a principled approach to select the most suitable neural network architecture for any given conjecture, and introduce a new dataset of graphs labeled with their Laplacian spectra. Furthermore, we provide a counterexample for a conjecture regarding the sum of the matching number and the spectral radius, which is simpler than the example provided in Wagner’s original paper.
The games have been implemented as environments in the Gymnasium framework, and along with the dataset, are available as open-source supplementary materials.
@@ 1 Introduction
11footnotetext: EU Horizon 2020: G.A. 871042 SoBig-Data++, NextGenEU - PNRR-PEAI (M4C2, investment 1.3) FAIR and “SoBigData.it”.22footnotetext: Funded by INdAM groups GNAMPA and GNSAGA.
@@
In this very general framework, once a target conjecture is chosen, say $f(G)<0$ for every graph $G$, there are several pivotal choices that could lead to success or failure. For instance, the rules of the “build your graph” game; the reward function; the termination condition; the RL algorithm used to play the game and optimize the cumulated reward; the neural network architecture involved in the RL algorithm. Each of these variables, and many others as well, has an impact on the model’s capability to explore successfully the space of graphs and eventually finding a counterexample. Moreover, if the bound is not surpassed despite the player learning, something can still be inferred: the conjecture is true, and experiments gives us empirical evidence in favor of this, or the counterexample is rare with respect to the visitation distribution of the RL algorithm in use, and this suggests changing some of the choices toward a more sophisticated exploration.
Novel Contributions. The aim of this paper is to open a discussion about those pivotal choices: among various “build your graph” games, reward functions, RL algorithms, neural network architectures, are some better than others? We argue that the first and most effective choice is the game, and we provide open-source Gymnasium implementations of four different graph-building games, that we call Linear, Local, Global, and Flip. An externally defined reward function makes them independent from the conjecture. Moreover, we argue that the second important choice is the neural network architecture, that should excel at extracting features informative for computing $f$. We recommend preliminary testing of various architectures on a supervised task related to $f$, selecting the one that performs best. We introduce a novel dataset of graphs labeled with their Laplacian spectra, which is particularly useful for conjectures related to eigenvalues. Furthermore, we present a novel counterexample for Conjecture 2.1 in [10]. With this contribution, we hope to steer the research toward a general systematization of Wagner’s framework.
@@ 2 Related Work
Wagner’s approach has led to several important follow-up studies, including [8], co-authored by Wagner himself. This study tackles an extremal graph theory problem originally proposed by Erdős in 1975. Their focus is on identifying graphs of a specific size that maximize the number of edges while excluding 3- or 4-cycles. Utilizing AlphaZero, they bootstrap the search process for larger graphs using optimal solutions derived from smaller ones, enhancing lower bounds across various sizes. Key innovations of their work include a new game, that they call edge-flipping game, and a novel Graph Neural Network architecture, called pairformer, which has proven particularly effective for this problem. However, they did not make the code for the game, the pairformer, or the AlphaZero configuration used in their experiments publicly available. In our paper, we provide Flip, a Gymnasium implementation of their edge-flipping game.
Another very interesting paper is [6], in which the authors reevaluate Wagner’s method with an emphasis on enhancing its speed and stability. They reimplement from scratch Wagner’s code, improving the readability, stability, and speed. They also successfully construct counterexamples for various conjectured bounds on the Laplacian spectral radius of graphs. Like our work, they implement an external reward function. Yet, the most important contribution of their paper is, in our opinion, the special attention given to computational performance. They observe that, since RL must process invariants for hundreds of thousands of graphs to achieve adequate convergence in learning, using networkX and/or numpy is suboptimal. They show that invoking Java code directly from Python significantly accelerates the invariants computation. Our experiments confirm that using networkX is quite slow, and using this approach is the most natural future development of our paper, see Section 6.
In  [7], authored by the same team as  [6], new lower bounds are established for several small Ramsey numbers. They continue to use Wagner’s original framework, but they with a slight modification to the RL algorithm. This aligns with our proposal to diversify Wagner’s framework from multiple directions to effectively tackle various conjectures.
@@ 3.1 Notation
In our environments, $\mathcal{G}$ is the family of all undirected graphs with a fixed amount $n$ of nodes, without multiple edges. Self-loops are optional.
@@ 3.2 States, Actions, and Transitions
In Wagner’s framework, the agent visits one edge at time, under a predefined order. However, in general, there are several different ways in which the agent can move on the graph and select the edge (or the edges) to modify. For instance, some games can be single-action based, like the edge-flipping game in [8] and our Flip game; other games can be multi-action, with both possibilities to modify an edge or not, like Wagner’s game, and our Linear, Local, and Global games. Moreover, the agent can modify an edge in several ways: by adding or removing it, but also by leaving as it is or changing it, or by flipping it. A significant aspect of how actions are defined is their impact on the game’s dynamics. In some cases, the game becomes monotonic, meaning that each graph $G_{t}$ is a subgraph of $G_{t+1}$. For details on our implementation and how we take these variations into account, see Section 4.
@@ 3.3 Reward
Given a conjecture as in (1), a natural choice for the reward in the episodic setting is $f(G_{T})$ at the end $T$ of the episode, and zero elsewhere. We call this reward sparse. Then, a counterexample is found when $r(G_{T})>0$ happens. A different choice is what we call the incremental reward: at each time $t>0$, we receive the increment $f(G_{t})-f(G_{t-1})$. If we start from a “virtual” graph $G_{0}$ with $f(G_{0})=0$, then the cumulated incremental reward is exactly $f(G_{T})$, as with the sparse reward. In certain cases, for instance with temporal difference algorithms, or when we are interested in understanding how a single action impacts on the graph, an incremental reward could prove useful.
Observe that in a graph-building game, reaching the end of the episode before checking for the counterexample is not efficient. Since the game is just a way to guide agent’s search for a counterexample, one could perform the check after every action. In this case, again, a sparse reward seems less reasonable. We provide the option of performing this check after every action in our envs. Note that in particular when this check is enabled, performing the invariant computation with an external Java code as suggested in [6], is particularly useful. See Section 6.
@@ 3.4 RL algorithm
The choice of reinforcement learning (RL) algorithm is crucial. Wagner, and [6, 7] as well, showed that even a simple, gradient-free algorithm like the cross-entropy method can achieve impressive results. However, there are several other algorithms that could be used. Among the most promising are Proximal Policy Optimization (PPO), which is known for its robustness and effectiveness across various tasks, and AlphaZero-like algorithms. Using PPO with our environments enabled us to find a novel counterexample with 18 nodes to Conjecture 2.1 in Wagner’s paper, see Figure 1.
Despite PPO being promising, we posit that AlphaZero is the most natural choice in this highly complex graph-building game scenario, and in fact it was used successfully in [8]. However, notice that while PPO is available in several established RL library like Stable-Baselines, and its design space is relatively easy to configure, AlphaZero complexity requires a much more thorough calibration of its hyperparameters.
@@ 3.5 Architecture of approximators
Given the combinatorial explosion of non-isomorphic graphs when increasing the number of nodes, approximation must be used. We assume that approximation is done by neural networks, and this takes the neural network architecture in play. Given the nature of the task, a clear “best choice” here is using a Graph Neural Network (GNN). But which GNN is best for a given conjecture $f$?
@@ 4 Environments
We implemented four “build your graph” games as Gymnasium environments: Linear, Local, Global, and Flip. All our games are played on undirected graphs with a fixed number $n$ of nodes, and without multiple edges. The environments are parametric with respect to several aspects: for instance, one can choose the starting graph, or whether to enforce the agent excluding self-loops. For details on these parameters, see Section 4.5
@@ 4.1 Linear
Linear is a variation of the game used by Wagner. The name comes from the state’s vector internal representation. In Linear, edges are ordered, and then at each time $t$ the agent can choose between leaving the edge number $t$ as it is, or flipping it. The state is given by the graph and the current time $t$, and the action space is $\{0,1\}$, where 00 means that the current edge is left unchanged, and $1$ that the current edge is flipped. With its default values, Linear differs from Wagner’s game with respect to the ordering of the edges: in Wagner’s game, edges are numbered by forming and expanding cliques first, that is, $(1,2),(1,3),(2,3),(1,4),\dots$, while in Linear the order is given by $(1,2),(1,3),\dots,(1,n),(2,3),\dots$ Moreover, Wagner starts from the empty graph, while the default setting in our games is to start from the complete graph. Episodes in Linear always end at time $T=\frac{n(n-1)}{2}$, if self-loops are not allowed, and at time $T=\frac{n(n+1)}{2}$, otherwise.
@@ 4.2 Local
In Local, the agent explores the graph space by moving from one node to another. When moving from node $i$ to node $j$, the agent has the option either to flip the edge $(i,j)$ or to take no action. This ensures that from node $i$, the agent’s actions are “locally” confined, impacting only the directly connected edge $(i,j)$. Note that this is different from Linear, because the agent can choose any node $j$ to move to. The state is given by the current graph, the current node $i$ where the agent is located, and the current time. An action is given by a target node $j$ to move to from node $i$, and a binary value $\{0,1\}$, where 00 means taking no action, and $1$ means flipping the edge $(i,j)$. Episodes in Local end at a termination time $T$ that can be passed as optional input when the game is initialized, and defaults to $T=\frac{n(n-1)}{2}$, if self-loops are not allowed, and to $T=\frac{n(n+1)}{2}$, otherwise.
@@ 4.3 Global
In Global, the agent explores the graph space by acting on any edge across the entire graph at any time. The agent can choose any edge to act upon, deciding either to flip it or to take no action. This “global” approach ensures that the agent’s actions are not confined to its immediate location, allowing interaction with any part of the graph. Episodes end at a termination time $T$ that can be given as input at game’s initialization, with same Local defaults. Similar to Local, the possibility to pass on an action without flipping an edge is maintained, because it helps mitigate the risk of choosing a wrong termination time for the game. For instance, if flipping were mandatory, excessively long matches could potentially disrupt an optimal configuration previously achieved. Allowing the passing of actions enables the agent to maintain an optimal configuration indefinitely. The state is given by the current graph, and the current time. An action is given by a target edge, and a binary value $\{0,1\}$, where 00 means taking no action, and $1$ means flipping the edge.
@@ 5 Dataset for Laplacian spectrum supervised learning
  • 1000 graphs drawn from Erdős–Rényi (ER) models, denoted as $G(11,p)$. We considered $p$ varying in $[0,1]$ with step $0.01$, obtaining 100 non-trivial models, and two trivial distributions that create copies of the empty graph or the complete graph. We have drawn $10$ graphs from each of this models. Our choice to produce a small samples size at fixed $p$ is motivated by the fact that slightly similar probabilities $p_{1}$ and $p_{2}$ lead to similar ER models, and thus, to similar graphs. Small models’ sample sizes produces a limited redundancy of configurations, resulting in a oversampling effect that helps the learning process.
@@
  • 8694 graphs obtained with Barabási–Albert model (BA), with parameter $m$ varying in $\{2,\dots,9\}$. BA algorithm builds a graph starting from an initial configuration on $m_{0}>m$ nodes. We took House of Graphs’ samples with $3\leq n\leq 10$ and used each $G$ in this batch to start a BA generation with $m<|V(G)|$.
The dataset is open-sourced and accessible in the supplementary material under the filename n11_graphs.g6. The .g6 format is a compact text-based encoding specifically designed for graphs. It is well-supported and can be easily read by the NetworkX method read_graph6, that returns a list of graphs objects. Labels are stored in the supplementary material under the filename n11_laplacian_spectra.txt. This is a simple text file, where each line includes 11 Laplacian eigenvalues in descending order, separated by spaces. Each line in the n11_laplacian_spectra.txt directly corresponds to the graph at the same position in the n11_graphs.g6 file.

### v2

@@ Abstract
In 2021, Adam Zsolt Wagner proposed an approach to disprove conjectures in graph theory using Reinforcement Learning (RL). Wagner frames a conjecture as $f(G)<0$ for every graph $G$, for a certain invariant $f$; one can then play a single-player graph-building game, where at each turn the player decides whether to add an edge or not. The game ends when all edges have been considered, resulting in a certain graph $G_{T}$, and $f(G_{T})$ is the final score of the game; RL is then used to maximize this score. This brilliant idea is as simple as innovative, and it lends itself to systematic generalization. Several different single-player graph-building games can be employed, along with various RL algorithms. Moreover, RL maximizes the cumulative reward, allowing for step-by-step rewards instead of a single final score, provided the final cumulative reward represents the quantity of interest $f(G_{T})$. In this paper, we discuss these and various other choices that can be significant in Wagner’s framework. As a contribution to this systematization, we present four distinct single-player graph-building games. Each game employs both a step-by-step reward system and a single final score. We also propose a principled approach to select the most suitable neural network architecture for any given conjecture and introduce a new dataset of graphs labeled with their Laplacian spectra. The games have been implemented as environments in the Gymnasium framework, and along with the dataset and a simple interface to play with the environments, are available at https://github.com/CuriosAI/graph_conjectures.
Keywords:
Reinforcement Learning Graph Theory.
77footnotetext: EU Horizon 2020: G.A. 871042 SoBig-Data++, NextGenEU - PNRR-PEAI (M4C2, investment 1.3) FAIR and “SoBigData.it”.88footnotetext: PRIN project Grafia (CUP: E53D23005530006), Department of Excellence MatMod@Tov (CUP: E83C23000330006).99footnotetext: Funded by INdAM groups GNAMPA and GNSAGA.footnotetext: ${}^{\text{all}}$omputational resources provided by CLAI laboratory, Chieti-Pescara, Italy.footnotetext: ${}^{\text{all}}$uthors can be contacted at curiosailab@gmail.com.
@@ 1 Introduction
In this very general framework, once a target conjecture is chosen, e.g. $f(G)<0$ for every graph $G$, there are several pivotal choices that could lead to success or failure. For instance, the rules of the “build your graph” game; the reward function; the termination condition; the RL algorithm used to play the game and optimize the cumulated reward; the neural network architecture involved in the RL algorithm. Each of these variables, and many others as well, has an impact on the model’s capability to explore successfully the space of graphs and eventually finding a counterexample. Moreover, if the bound is not surpassed despite the player learning, something can still be inferred: the conjecture is true, and experiments gives us empirical evidence in favor of this, or the counterexample is rare with respect to the visitation distribution of the RL algorithm in use, and this suggests changing some of the choices toward a more sophisticated exploration.
Novel Contributions. The aim of this paper is to open a discussion about those pivotal choices: among various “build your graph” games, reward functions, RL algorithms, neural network architectures, are some better than others? We argue that the first and most effective choice is the game, and we provide open-source Gymnasium [18] implementations of four different graph-building games, that we call Linear, Local, Global, and Flip. An externally defined reward function makes them independent from the conjecture. Moreover, we argue that the second important choice is the neural network architecture, that should excel at extracting features informative for computing $f$. We recommend preliminary testing of various architectures on a supervised task related to $f$, selecting the one that performs best. We introduce a novel dataset of graphs labeled with their Laplacian spectra, which is particularly useful for conjectures related to eigenvalues. Furthermore, we present a novel counterexample for Conjecture 2.1 in [19]. With this contribution, we hope to steer the research toward a general systematization of Wagner’s framework.
@@ 2 Related Work
Wagner’s approach has led to several important follow-up studies, including [9], co-authored by Wagner himself. This study tackles an extremal graph theory problem originally proposed by Erdős in 1975. Their focus is on identifying graphs of a specific size that maximize the number of edges while excluding 3- or 4-cycles. Utilizing AlphaZero[15], they bootstrap the search process for larger graphs using optimal solutions derived from smaller ones, enhancing lower bounds across various sizes. Key innovations of their work include a new game, that they call edge-flipping game, and a novel Graph Neural Network architecture, called pairformer, which has proven particularly effective for this problem. However, they did not make the code for the game, the pairformer, or the AlphaZero configuration used in their experiments publicly available. In our paper, we provide Flip, a Gymnasium implementation of their edge-flipping game.
Another very interesting paper is [6], in which the authors reevaluate Wagner’s method with an emphasis on enhancing its speed and stability. They reimplement from scratch Wagner’s code, improving the readability, stability, and speed. They also successfully construct counterexamples for various conjectured bounds on the Laplacian spectral radius of graphs. Like our work, they implement an external reward function. Yet, the most important contribution of their paper is, in our opinion, the special attention given to computational performance. They observe that, since RL must process invariants for hundreds of thousands of graphs to achieve adequate convergence in learning, using NetworkX [8] and/or numpy is suboptimal. They show that invoking Java code directly from Python significantly accelerates the invariants computation. Our experiments confirm that using networkX is quite slow, and using this approach is the most natural future development of our paper, see Section 6.
In  [7], authored by the same team as  [6], new lower bounds are established for several small Ramsey numbers. They continue to use Wagner’s original framework, but with a slight modification to the RL algorithm. This aligns with our proposal to diversify Wagner’s framework from multiple directions to effectively tackle various conjectures.
@@ 3.1 Notation
In our environments, $\mathcal{G}$ is the family of all undirected unweighted graphs with a fixed amount $n$ of nodes, without multiple edges. Self-loops are optional.
@@ 3.2 States, Actions, and Transitions
In Wagner’s framework, the agent visits one edge a time, under a predefined order. However, in general, there are several different ways in which the agent can move on the graph and select the edge (or the edges) to modify. For instance, some games can be single-action based, like the edge-flipping game in [9] and our Flip game; other games can be multi-action, with both possibilities to modify an edge or not, like Wagner’s game, and our Linear, Local, and Global games. Moreover, the agent can modify an edge in several ways: by adding or removing it, but also by leaving it as it is or changing it, or by flipping it. A significant aspect of how actions are defined is their impact on the game’s dynamics. In some cases, the game becomes monotonic, meaning that each graph $G_{t}$ is a subgraph of $G_{t+1}$. For details on our implementation and how we take these variations into account, see Section 4.
@@ 3.3 Reward
Given a conjecture as in (1), a natural choice for the reward in the episodic setting is $r(G_{T})=f(G_{T})$ at the end $T$ of the episode, and $r(G_{T})=0$ elsewhere. We call this reward sparse. Then, a counterexample is found when $r(G_{T})>0$ happens. A different choice is what we call the incremental reward: at each time $t>0$, we receive the increment $f(G_{t})-f(G_{t-1})$. If we start from a “virtual” graph $G_{0}$ with $f(G_{0})=0$, then the cumulated incremental reward is exactly $f(G_{T})$, as with the sparse reward. In certain cases, for instance with temporal difference algorithms, or when we are interested in understanding how a single action impacts on the graph, an incremental reward could prove useful.
Observe that in a graph-building game, reaching the end of the episode before checking for the counterexample is not efficient. Since the game is just a way to guide agent’s search for a counterexample, one could perform the check after every action. In this case, again, a sparse reward seems less reasonable. We provide the option of performing this check after every action in our environments. Note that in particular when this check is enabled, performing the invariant computation with an external Java code as suggested in [6], is particularly useful. See Section 6.
@@ 3.4 RL algorithm
The choice of the reinforcement learning (RL) algorithm is crucial. Wagner, and [6, 7] also showed that even a simple, gradient-free algorithm like the cross-entropy method can achieve impressive results. Nonetheless, several alternative algorithms could also be considered. Among the most promising are Proximal Policy Optimization (PPO), known for its robustness and effectiveness across various tasks, and AlphaZero-like algorithms. Employing PPO with our environments enabled us to find a novel 18-node counterexample to Conjecture 2.1 in Wagner’s paper, see Fig. 1. Although this counterexample merely replicates the structure of Wagner’s original, it underscores the potential applicability of different algorithms.
Despite PPO being promising, we posit that AlphaZero is the most natural choice in this highly complex graph-building game scenario, and in fact it was used successfully in [9]. However, notice that while PPO is available in several established RL library like Stable-Baselines [14], and its design space is relatively easy to configure, AlphaZero complexity requires a much more thorough calibration of its hyperparameters.
@@ 3.5 Architecture of approximators
Given the combinatorial explosion of non-isomorphic graphs when increasing the number of nodes, approximation must be used. We assume that approximation is done by neural networks, and this takes the neural network architecture in play. Given the nature of the task, a clear “best choice” here is using a Graph Neural Network (GNN). But which GNN is the best one for a given conjecture $f$?
@@ 4 Environments
We implemented four “build your graph” games as Gymnasium environments: Linear, Local, Global, and Flip. All our games are played on undirected graphs with a fixed number $n$ of nodes, and without multiple edges. The environments are parametric with respect to several aspects: for instance, one can choose the starting graph, or whether to enforce the agent excluding self-loops. For details on these parameters, see Section 4.5. These environments are available at https://github.com/CuriosAI/graph_conjectures.
@@ 4.1 Linear
Linear is a variation of the game used by Wagner. The name comes from the state’s vector internal representation. In Linear, edges are ordered, and then at each time $t$ the agent can choose between leaving the edge number $t$ as it is (i.e. passing it), or flipping it. The Edge-flipping operation (as defined in [9]) changes the state of an edge like a boolean not operator, as follows: let $e\in\{0,1\}$ be the single bit representing the edge, then
$flip(e)=\begin{cases}1\quad&\text{if }e=0\\ 0&\text{otherwise}\end{cases}$
The state is given by the graph and the current time $t$, and the action space is $\{0,1\}$, where 00 means that the current edge is left unchanged, and $1$ that the edge is flipped. With its default values, Linear differs from Wagner’s game for the ordering of the edges: in Wagner’s game, edges are numbered by forming and expanding cliques first, that is, $(1,2),(1,3),(2,3),(1,4),\dots$, while in Linear is given by $(1,2),(1,3),\dots,(1,n),(2,3),\dots$. Moreover, Wagner starts from the empty graph, while the default setting in our games is to start from the complete graph. Episodes in Linear always end at time $T=\frac{n(n-1)}{2}$, if self-loops are not allowed, and at time $T=\frac{n(n+1)}{2}$, otherwise.
G123412341243
Figure 2: Effects of the flip action on different edges. Existing and missing edges are represented with solid and dotted lines, respectively. The top-right graph shows G after flipping edge (2,3), while the bottom-right graph shows G after flipping edge (1,4).
@@ 4.2 Local
In Local, the agent explores the graph space by moving from one node to another. When moving from node $i$ to node $j$, the agent has the option either to flip the edge $(i,j)$ or to pass it. This ensures that from node $i$, the agent’s actions are “locally” confined, impacting only the directly connected edge $(i,j)$. Note that this is different from Linear, because the agent can choose any node $j$ to move to. The state is given by the current graph, the current node $i$ where the agent is located, and the current time. An action is given by a target node $j$ to move to from node $i$, and a binary value $\{0,1\}$, where 00 means taking no action, and $1$ means flipping the edge $(i,j)$. In our implementation, this action logic is represented by a single integer value $k$ within the range $[0,2n-1]$ where $n$ is the number of nodes. Assuming to start from node $i$, if $k\in[0,n-1]$, we move to node $j=k$ without taking any action on edge $(i,j)$. If $k\in[n,2n-1]$, we move to node $j=k\mod n$, and the edge $(i,j)$ is flipped. Episodes in Local end at a termination time $T$ that can be passed as optional input when the game is initialized, and defaults to $T=\frac{n(n-1)}{2}$, if self-loops are not allowed, and to $T=\frac{n(n+1)}{2}$, otherwise.
@@ 4.3 Global
In Global, the agent explores the graph space by acting on any edge across the entire graph at any time. The agent can choose any edge to act upon, deciding either to flip it or to pass it. This “global” approach ensures that the agent’s actions are not confined to its immediate location, allowing interaction with any part of the graph. Episodes end at a termination time $T$ that can be given as input at game’s initialization, with same Local defaults. Similar to Local, the possibility to pass on an action without flipping an edge is maintained, because it helps mitigate the risk of choosing a wrong termination time for the game. For instance, if flipping were mandatory, excessively long matches could potentially disrupt an optimal configuration previously achieved. Allowing the passing of actions enables the agent to maintain an optimal configuration indefinitely. The state is given by the current graph, and the current time. An action is given by a target edge, and a binary value $\{0,1\}$, where 00 means taking no action, and $1$ means flipping the edge. In our implementation, the action logic is similar to that seen in Local, but generalized to handle global movements along the graph. Here, the action is represented by a single integer value $k$ within the range $[0,2m-1]$, where $m$ is the number of edges. If $k\in[0,m-1]$, edge $(i,j)$, where $i=\left\lfloor\frac{k}{n}\right\rfloor$ and $j=k\mod m$, remains unchanged. If $k\in[m,2m-1]$, edge $(i,j)$, where $i=\left\lfloor\frac{k-m}{n}\right\rfloor$ and $j=k\mod m$ is flipped.
@@ 4.4 Flip
(a) Linear
(b) Local
(c) Global and Flip
Figure 3: Comparison of game modes. Edges that can be modified by the agent are colored blue, with dotted lines representing missing edges. In Linear the agent can modify just the next edge in the game’s order. In Local, the agent remains on vertices, such as the blue one in 3(b), and can choose to modify one among its incident edges. In Global and Flip, all edges are always accessible.
@@ 4.5 Common settings
4.6 User interface
To facilitate testing of the game environments described herein, we provide a simple graphical user interface (GUI), which is open-sourced and available at https://github.com/CuriosAI/graph_conjectures under main.py. This interface enables users to select a conjecture (Wagner or Brouwer), choose a game type (Linear, Local, Global, or Flip), specify the number of nodes, and select the reward type (Sparse or Incremental). Users can interact with the game by performing actions, with the GUI visualizing the current state of the game as graph $G$, as well as the function $f(G)$ that is intended to be maximized.
@@ 5 Dataset for Laplacian spectrum supervised learning
  • 1010 graphs drawn from Erdős–Rényi (ER) models, denoted as $G(11,p)$. We considered $p$ varying in $[0,1]$ with step $0.01$, obtaining 99 non-trivial models, and two trivial distributions that create copies of the empty graph or the complete graph. We have drawn $10$ graphs from each of this models. Our choice to produce a small samples size at fixed $p$ is motivated by the fact that slightly similar probabilities $p_{1}$ and $p_{2}$ lead to similar ER models, and thus, to similar graphs. Small models’ sample sizes produces a limited redundancy of configurations, resulting in a oversampling effect that helps the learning process.
@@
  • 10162 graphs obtained with Barabási–Albert model (BA), with parameter $m$ varying in $\{2,\dots,9\}$. BA algorithm builds a graph starting from an initial configuration on $m_{0}>m$ nodes. We took House of Graphs’ samples with $3\leq n\leq 10$ and used each $G$ in this batch to start a BA generation with $m<|V(G)|$.
The dataset is open-sourced and available at https://github.com/CuriosAI/graph_conjectures under the filename n11_graphs.g6. The .g6 format is a compact text-based encoding specifically designed for graphs. It is well-supported and can be easily read by the NetworkX method read_graph6, that returns a list of graphs objects. Labels are in n11_laplacian_spectra.txt. This is a simple text file, where each line includes 11 Laplacian eigenvalues in descending order, separated by spaces. Each line in the n11_laplacian_spectra.txt directly corresponds to the graph at the same position in the n11_graphs.g6 file.
All pairs of graphs in the dataset have been subjected to the 1-dimensional Weisfeiler-Leman test [20]. This test gives a negative response in case of non-isomorphic graphs, and positive in case of potentially isomorphic graphs (false positives are possible). The test resulted positive on 1124695 pairs, meaning that the percentage of isomorphic pairs is less than or equal to the 1,57% of the total. The highlighted pairs are reported in the file weisfeiler_leman_results.txt.